### v2

# An Empirical Study of Cross-Lingual Transfer Learning in Programming Languages

Abstract
Large language models (LLMs) have become remarkably good at improving developer productivity for high-resource programming languages. These models use two kinds of data: large amounts of unlabeled code samples for pre-training and relatively smaller amounts of labeled code samples for fine-tuning or in-context learning. Unfortunately, many programming languages are low-resource, lacking labeled samples for most tasks and often even lacking unlabeled samples. Therefore, users of low-resource languages (e.g., legacy or new languages) miss out on the benefits of LLMs. Cross-lingual transfer uses data from a source language to improve model performance on a target language. It has been well-studied for natural languages, but has received little attention for programming languages. This paper reports extensive experiments on four tasks using a transformer-based LLM and 11 to 41 programming languages to explore the following questions. First, how well does cross-lingual transfer work for a given task across different language pairs. Second, given a task and target language, how should one choose a source language. Third, which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task. Our empirical study with 1,808 experiments reveals practical and scientific insights, such as Kotlin and JavaScript being the most transferable source languages and different tasks relying on substantially different features. Overall, we find that learning transfers well across several programming languages.
Figure 1: Overview. Consider an LLM pretrained on unlabeled code in multiple seen languages. Finetune on task-specific labeled samples from a source language. For RQ1, test performance on a target language. For RQ2, train a model that ranks transfer performance given features of a language pair. For RQ3, measure how important the features of language pairs are. Repeat for several languages and four tasks: Error Detection, Solution Domain Classification, Clone Detection, and Code Repair.
I Introduction
Large language models (LLMs) leverage the naturalness of software [1] to achieve state-of-the-art performance on many software engineering tasks such as error detection, clone detection, and code translation [2, 3]. However, LLMs have thus far been trained and evaluated mainly on programming languages with vast amounts of openly available code rather than low-resource ones. Programming languages used for critical and strategic applications may have much code only in private repositories and are thus ignored. COBOL, which supports many critical financial applications and may have over 775 billion lines of code overall [4], is low-resource due to lack of open availability. Similarly with Fortran, which is used in high performance and scientific computing domains. Notwithstanding their popularity [5, 6], new languages such as Rust and Swift start out low-resource, by definition. Other languages such as R may be low-resource for licensing reasons [7]. Besides being underrepresented in training of LLMs, organizations often spend significant resources on training software developers to work with unfamiliar low-resource languages. This adds to the cost of maintaining such software and may also stifle innovation [8]. There is a need for state-of-the-art AI tools to augment the capabilities of developers for these low-resource languages. Indeed, LLMs are being used in tools to enhance developer productivity and can also help with migration or modernization of projects from one programming language to another by supporting code translation [9, 10].
Recent works have shown the possibility of transfer learning: leveraging data from one programming language to compensate for the lack of data in the target programming language [11, 12, 13]. Another recent work has developed several similarity metrics to decide which high-resource programming language dataset can be used to augment fine-tuning data for some other language [14]. Prior work, however, has used at most 6 programming languages, none of which are truly low-resource. Further, the notion of similarity among programming languages is underexamined: more research is needed to make reliable claims. To overcome these shortcomings, we examine transferability from source to target language combinations for 4 tasks and for many more languages than prior work (11–41 depending on the task). The four tasks we focus on are clone detection, code repair, solution domain classification, and error detection. Fig. 1 gives a high-level overview of the methodology for our extensive empirical study. From left to right, we start with a pretrained LLM that has seen unlabeled code in several programming languages. Next, for each of four tasks, for each source language with a suitable amount of labeled training data for that task, we finetune the LLM. For each target language with a suitable amount of labeled test data for the task, we evaluate the finetuned LLM. For example, Fig. 1 illustrates this with a solution domain classification task, using C as the source language and Rust as the target language, which yields an F1 score of $0.48$. Doing this for four different tasks and all their source and target language combinations yields four different heat maps.
Our first research question (RQ1) asks how learning transfers across programming languages in general, by exploring these heat maps. Our second research question (RQ2) asks whether one can predict the ranking of source languages in the heat maps from the previous experiment just using features of language pairs, by training a performance prediction model from the language pair features and the ground-truth target labels from the heat maps. For example, in Fig. 1, the performance prediction model for the solution domain classification task might rank C behind Kotlin as a source language for transfer to Rust. Given the cost of training LLMs, the performance prediction model is useful in directing data acquisition efforts and deciding how to spend compute resources. Finally, as our third research question (RQ3), we seek software-engineering insights into the characteristics of programming languages that affect transferability. To do this, we measure the importance of language pair features in the prediction model from the previous experiment.
This paper makes the following contributions:
  • Evaluating the pairwise transferability for several programming languages (including low-resource ones) and four tasks using LLMs.
  • Developing a method to identify the best language to transfer from, for different target languages and tasks.
  • Characterizing the features of programming language pairs that are predictive of transferability for given tasks.
We are planning to make our data generation, training, and analysis code available. One goal of this paper is to offer practical guidance to engineers who build LLM-driven software engineering tools towards more informed choices for data acquisition and modeling. A second goal of this paper is to help advance software engineering as a science by using the lens of transfer learning to shed light on programming language characteristics. Below are some results highlights:
  • Learning transfers well for all tasks. More specifically, cross-lingual learning transfers better than zero-shot.
  • Transfer learning depends on the target language. Java, Go, Dart, and TypeScript are among the best target languages. C++, Python, Vim, and COBOL are among the worst.
  • Transfer learning also depends on source language used for finetuning. Kotlin and JavaScript are the best source languages, C++ is the worst.
  • We can predict which source language can perform well for a given task by considering a variety of features.
  • Different tasks rely on substantially different features.
  • Overall the most important features are the overlap of keywords and names between the source and target languages.
If we take the premise that software engineers will increasingly benefit from LLM support, then we hope this paper will help democratize such support into low-resource settings.
II Background and Related Work
Cross-lingual Transfer for Natural Languages
Lin et al. [15] use 4 tasks to study transfer among up to 60 natural languages and explore how important features of language pairs are for predicting how well learning transfers between them, finding that feature importances vary a lot across tasks. Lauscher et al. [16] use 5 tasks to study transfer among up to 15 languages and argue that while zero-shot performance can work for low-level tasks, higher-level tasks benefit from at least a few target-language shots. de Vries et al. [17] use 1 task (POS tagging) with 65 source and 105 target languages and study the effects of language families and writing systems, with Romanian emerging as a particularly good source language. Ahuja et al. [18] use 11 tasks with 1 source and varying numbers of target languages, confirming earlier findings that feature importances vary a lot across tasks. Our paper takes inspiration from these works, especially de Vries et al. [17], but differs by studying programming languages instead of natural languages, which have different tasks and different features of language pairs affecting transferability.
Cross-lingual Transfer for Programming Languages
Zhou et al. [5] call for studying transfer, motivating with 1 task (code completion) and 2 languages (from Hack to Rust). Chen et al. [14] use 2 tasks (code summarization and search) and 6 languages to study transfer to Ruby, and propose picking the source language based on language similarity. Ahmed et al. [11] use 3 tasks to study transfer among 6 languages, demonstrating that due to the nature of their tasks, signals from identifiers are highly important for transferability. Yuan et al. [12] use 1 task (automated program repair) to study transfer among 5 languages, sequentially fine-tuning on multiple languages with innovative tricks to prevent catastrophic forgetting. Pian et al. [13] use 2 tasks (code summarization and completion) to study transfer among 4 languages, using meta-learning to improve a base learner. Our paper also focuses on programming languages, but considers more tasks and vastly more languages to explore conditions for effective transfer.
Besides the above works, there is also work on learning for multiple programming languages. While such prior work has been instrumental in making studies like ours possible, it differs by not focusing on cross-lingual transfer. Our work benefits from datasets for multiple programming languages. CodeNet covers 55 languages [19]; XCodeEval covers 7 tasks with up to 17 languages [20]; and MultiPL-E supports 1 task (code generation) in 19 languages [21]. There are also models for multiple programming languages. TransCoder uses pre-training across 3 languages to help transpilers [10]. CodeT5 is pre-trained on 8 languages and 5 tasks [22]. StarCoder is pretrained on 86 languages [7], but e.g. does not include COBOL and evaluation is largely focused on Python. Unsupervised pretraining on multiple languages has become common, but transferability of supervised tasks has not yet been thoroughly studied; our paper addresses that gap.
III Experimental Setup
As illustrated in Fig. 1, given a task, our experimental approach first finetunes a model for each source language individually and then tests each finetuned model on all target languages. We applied this approach to four tasks where the number of source languages varies from 6 to 22, leading to 58 finetuned models in total. For each source language, we finetune the model using one A100 GPU for 6-30 hrs depending on the task. Each model was then evaluated on 11 to 43 target languages producing 1,808 experiments. All results are presented in Figs. 5(a)–(d) and analyzed in Sec. IV.
This section presents methodological details.
(a) Clone Detection: 21 source languages $\times$ 41 target languages. Metric: F1 Score.
Figure 2: Transfer scores heat map. The figure shows scores for every combination of source and target language. Each column corresponds to a source language, with “zsh” showing zero-shot performance, i.e., without fine-tuning on any source language. Each row corresponds to a target language. The languages whose language-name label uses red font were seen during pre-training. Framed black boxes highlight the performance of a source language (column) on itself as the target language (row). The dendrograms show results of hierarchical clustering based on similarity of the performance vectors. The row and column order is also determined by the same clustering.
(a) Code Repair: 20 source languages $\times$ 38 target languages. Metric: BLEU score.
Figure 3: Transfer scores heat map (cont.; see caption from 2(a))
(a) Solution Domain Classification: 11 source languages $\times$ 11 target languages. Metric: F1 score.
Figure 4: Transfer scores heat map (cont.; see caption from 2(a)).
(a) Error Detection: 6 source languages $\times$ 11 target languages. Metric: F1 Score.
Figure 5: Transfer scores heat map (cont.; see caption from 2(a)).
Base Datasets
The four tasks studied are derived from two multilingual code datasets: CodeNet [19] and XCodeEval [20]. CodeNet [19] consists of about 14M code samples for a total of 500M lines of code in 55 different programming languages, derived from submissions to online coding judge websites. The dataset comes with benchmarks for code classification and code similarity. XCodeEval [20] contains 25M samples in 11 programming languages from another different online coding judge website. This dataset comes with an execution-based evaluation framework and several different benchmarks: from classification (solution domain classification and error detection) to generation (program synthesis, automatic program repair, and code translation).
Tasks
To study how learning transfers between programming languages, we explore two type of tasks: classification (error detection, solution domain classification, clone detection) and generation (code repair). For the classification tasks, we follow Lewis et al. [23] by predicting labels from the vocabulary of class labels based on the final decoder hidden state. We use BLEU score [^1] We chose BLEU score because it enables experiments with the larger number of languages in CodeNet [19], and because other prominent benchmarks for the code repair task also adopt this metric [24]. for evaluating the generative task and F1 score for evaluating the classification tasks.
  • Error Detection: Given a code $C$ in language $L$, the task is binary classification of whether or not $C$ compiles (or can be loaded by an interpreter) without error.
  • Solution Domain Classification: Given a code $C$ in language $L$, the task is multi-label classification of $C$ into a set of tags corresponding to potential algorithmic techniques required to write the solution (e.g., sorting, graphs).
  • Clone Detection: Given two code samples $C_{1}$ and $C_{2}$ in language $L$, the task is binary classification of whether they are type-IV clones (semantically similar) [25]. To generate the dataset, we apply the following procedure to the CodeNet dataset. Given all combinations of solutions to all problems in language $L$, we identify positive samples (clones) as pairs of accepted solutions for the same problem and the others as negative examples. To balance the positive samples across problems, we ensure a ratio of 0.15 of positive samples across different languages.
  • Code Repair: Given a buggy code $C$ in language $L$, the task is to generate the corresponding repaired code. We generate the dataset by using the following procedure on CodeNet solutions: Each sample is modified by sequentially inserting, removing, or replacing tokens of different types for fixed ratios for different token types.
Data Sampling
Given variation in the size of datasets for different languages, we follow the sampling procedure of de Vries et al. [17] to select the number of training samples and identify potential source languages. We first finetune the model with datasets of different number of samples $N$={10K, 30K, 50K, 70K, 100K}. Depending on the number of training examples $N_{L}$ for a language $L$, we randomly upsample languages with $N<N_{L}$ or downsample languages with $N>N_{L}$. We filter the set of source languages for a given task based on the relative performance compared to a baseline model. We select sample sizes of 50K for solution domain classification, 70K for error detection, and 100K for clone detection and code repair. Fig. 8 in the supplementary material shows how the metric for each task vaires with the sample size.
Model
Our experimental framework is based on CodeT5-base (220M parameters) leveraging the Hugging Face transformers library [26]. Code-T5 is an open-source model, pre-trained on eight programming languages (Ruby, JavaScript, Go, Python, Java, PHP, C, and C#). Due to its encoder-decoder nature, the model performs well on both code generation and code understanding tasks [22] like code repair, error detection, and clone detection, among others. This, along with its relatively small size, makes it a good fit for our empirical study, which requires 58 finetuned models with numerous inference runs each on four tasks.
IV Results and Discussion
This section addresses the following research questions.
RQ1:
How well does cross-lingual transfer work for a given task across different language pairs?
RQ2:
Given a task and target language, how should we pick the source language for best performance?
RQ3:
Which characteristics of a language pair are predictive of transfer performance, and how does that depend on the given task?
IV-A Transfer analysis
We explore RQ1 via extensive experiments on every combination of source and target programming language, for each task. The four heat maps of Fig. 5 show the results. A source programming language is the language used for fine-tuning a model, with samples in the training data. A target programming language is the language used to evaluate a model, with samples in the test data. Fig. 5 shows source languages on the X-axis and target languages on the Y-axis. Monolingual scores (where the source and the target language are the same) are highlighted with bold boxes.
For RQ1 we analyze model performance trends with respect to task, source language, and target language. We provide some practical insights, like which languages are the toughest target languages and which are the best source languages. We also discuss some interesting source-and-target language pairs.
Task dependency
The heat maps in Fig. 5 show that transferability of source languages varies depending on the task. Visually this can be seen by how the color gradient varies for each task. To quantify this observation, we calculate the mean scores across all the language combinations. The mean cross-lingual scores are 0.78 for clone detection, 0.75 for code repair, 0.65 for error detection, and 0.47 for solution domain classification. The monolingual scores (in bold boxes) are higher (more red) than the cross-lingual scores, averaging 0.91 for clone detection, 0.98 for code repair, 0.78 for error detection, and 0.51 for solution domain classification. The zero-shot columns, where the model is not finetuned on any source language, have lower scores (more blue) for all heat maps. The mean zero-shot scores are 0.49 for clone detection, 0.28 for code repair, 0.49 for error detection, and 0.01 for solution domain classification. Variations in the mean monolingual, cross-lingual, and zero-shot scores across tasks confirm that model performance is task dependent. The difference in overall zero-shot and cross-lingual scores indicates that in the absence of fine-tuning data for any task in any programming language, it is better to finetune with some other language in that task than using a zero-shot setting. This cross-lingual advantage over the zero-shot setting occurs for every task.
Looking at these scores in absolute terms reveals that the choice of tasks also affects transferability. For solution domain classification, the difference in mean cross-lingual and monolingual scores is only 0.04, whereas for code repair, it is 0.23. The difference in mean cross-lingual and zero-shot scores for the same two tasks is 0.46 and 0.48. Although overall trends show learning transfers well in all tasks, cross-lingual performance in relation to monolingual performance is highly task dependent. Within each task, model performance depends on the choice of source and target programming language, which we explore next. Tab. II in the supplementary material shows the mean score by task.
Target language dependency
Target programming languages fall into two categories, high- and low-resource. In Fig. 5, high-resource target programming languages are represented on both axes. Low-resource target programming languages are represented only on the Y-axis. Models are finetuned only on high-resource programming languages and tested on both high- and low-resource programming languages.
To understand overall trends across all four tasks, we consider ten target programming languages that are common to all four tasks. Six (Kotlin, JavaScript, Java, Python, C, and C++) are high-resource in all four tasks. Four (Go, Rust, PHP, and C#) are low-resource in error detection. For each of these ten languages, for each of the four tasks, we calculate the mean cross-lingual score using all scores in the row of Fig. 5 for each language, excluding the monolingual score in the bold box and the zero-shot score in the extreme left. We exclude monolingual and zero-shot scores so as to understand transferability from different languages to the target language. Based on this mean we rank each of the 10 languages for each task, and then calculate the mean rank for each language.
We find Java, Go, Rust, JavaScript, and Kotlin are target languages that benefit most from transfer learning, while PHP, C#, C, C++, and Python benefit least, with Java, Go being the best and Python, C++ being the worst. C++ has a reputation as being one of the hardest programming languages to learn for human software developers, which may explain our finding that it is a difficult target for LLM transfer learning. Some target programming languages are good at monolingual performance, but much worse at cross-lingual performance; one example across multiple tasks is C++. In-language performance for C++ is 0.95, 0.99, and 0.46 for clone detection, code repair, and solution domain classification, whereas its mean cross-language performance is 0.79, 0.81, and 0.41 for the same tasks. Java is an example of more consistent in-language and cross-language transferability across tasks and hence can be considered relatively more transferable than C++. Java has in-language scores of 0.95, 0.99, and 0.45 for clone detection, code repair, and solution domain classification, and cross-language scores of 0.86, 0.84, and 0.42 for the same. All numbers are reported Tab. III in the supplementary material.
We do a similar analysis for low-resource target languages in clone detection and code repair, since these two tasks have a wider and more varied set of low-resource languages. Here we find Dart, TypeScript, Elixir, Lua, and Swift benefit most from cross-lingual training, with Dart being the best. Dart, TypeScript have average scores of 0.87 and 0.79, 0.82 and 0.83 for clone detection and code repair. COBOL, F#, Octave, bc, and Vim are least likely to benefit, with Vim being the worst. Vim has average scores of 0.69, 0.67 and COBOL has average scores of 0.72, 0.74 for Clone Detection and Code Repair. Tab. IV in the supplementary material has more information on mean rank and average score of low-resource target languages for clone detection and code repair.
For error detection, only a smaller set of languages can be considered low resource. We observe low variation in average scores across different sources with highest being Rust with 0.66 and lowest being C# with 0.64. For solution domain classification there are no low-resource target languages.
Our analysis shows that transferability depends on the choice of task and target language. We can see by comparing with zero-shot scores that all target programming languages benefit from cross-lingual training, however some languages benefit much more than others.
Most Transferable Source Language
The number of source languages vary for each of the four tasks. To understand the source language dependency and to identify the most transferrable source language, we start from languages all four tasks have in common (JavaScript, C, Kotlin, Java, C++, and Python). For each task, for each language, we calulate the mean cross-lingual score. As an example, consider the column for source language Python on the X-axis of Fig. 5(a). We calculate the mean over the scores except 0.94, which is the monolingual score with target language Python. Once we have the cross-lingual mean for each language, we rank the source languages based on their mean score for each task. Based on these rankings, we calculate the mean rank, for overall cross-lingual performance of source languages, across all target languages, across all tasks.
The most transferrable source languages from best to worst are Kotlin, JavaScript, Java, Python, C, and C++. It may be surprising that Kotlin is the best-performing source language, since it is rarely discussed in the context of training LLMs and is not part of the pre-training corpus for CodeT5 [22]. Prior work on transfer learning for natural languages [17] yielded a similar surprise, with Romanian as the best-performing source language. A possible explanation could be that Romanian and Kotlin both have strong roots in dominant languages of the past (Latin and Java), while sitting at a cross-roads that helped them absorb many influences. Detailed scores and rankings are in Tab. V in the supplementary material.
The above analysis considers all target languages, including high-resource target languages. In a more realistic setting, we can ignore high-resource target languages because the monolingual option is available for every high-resource language in every task. If we repeat the above analysis with only low-resource languages, we are left with a smaller set of target languages and only three tasks (all target languages for solution domain classification are high-resource). For the remaining three tasks and six common source languages, analysis shows that the most transferrable source languages are JavaScript, Java, Kotlin, C, Python, and C++, with JavaScript being the best and C++ being the worst. JavaScript and Java contribute the most data to CodeT5 pretraining; Kotlin and C++ are absent [22]. For details see Tab. VI in the supplementary material.
If we consider only clone detection and code repair, we get a much broader set of about 20 common source languages. Repeating the analysis for these two tasks for low-resource target languages only, we find once again that JavaScript is the most transferable and C++ is among the least. In this case the most transferable languages are JavaScript, Scala, and Haskell, while the least transferable are Fortran, AWK, C++, Lisp, and Pascal. More details are in Tab. VII in the supplementary material.
Language Pair Performance
The best model performance for each task is almost always in the monolingual setting. However, some cross-lingual pairs do stand out and show very good model performance, comparable to monolingual. Another reason to consider individual language pairs is that from the perspective of developers working with low-resource languages, it would be helpful to know which programming languages act as the best source language.
For clone detection in Fig. 5(a), notice that a model finetuned on Ruby does almost as well on Crystal, a language whose syntax is inspired by Ruby. Fig. 5(a-b) also shows that surprisingly, Haskell does very well on COBOL, much better than FORTRAN does on COBOL. From the same heat maps notice Lisp does very well on Scheme, which can be expected as Scheme is a dialect of Lisp. Similary Python does very well on Cython. Fig. 5(a-d) also show that C-category languages (C, C++, and C#) do well on each other. Details are in Tab. VIII in the supplementary material.
Dendrograms
Dendrograms for each task cluster programming languages with similar transfer learning performances. In the code repair task, syntaxically and semantically related languages are next to each other. For example, C is paired with C++ and is next to the pair C# and Java. JavaScript is paired with TypeScript, Python with Cython, Lua with Moon, Lisp with Scheme. On the other hand, relatedness of languages plays a smaller role in the dendrograms for other tasks. This might be because while the code repair task requires deep language understanding, other tasks depend more upon other features. For instance, Ahmed and Devanbu [11] have shown that for some tasks, identifiers are more important. Section IV-C will analyze the relationship between features of language pairs and their transfer performance.
1. Learning transfers well for all tasks. Cross-lingual learning transfers better than zero-shot. 2. Learning depends on the target language. Among low resource languages, Dart and TypeScript are the best while Vim and COBOL are among the worst. 3. Kotlin and JavaScript are among the best target languages, C++ is among the worst. 4. Some cross-lingual language pairs show particularly good results, such as Ruby and Crystal, Haskell and COBOL.
IV-B Performance Prediction
Feature Measure Description Linguistic Features Object oriented e Is the language object oriented? Type strength e Is the language strongly or weakly typed? Type checking e Is the language typed statically or dynamically? Type safety e Is the language type safe? Garbage collection e Does the language use garbage collector? Standardized e Is the language standardized by a committee? Expression of types e Are the types written explicitly? Paradigm o Paradigms supported (e.g., functional, imperative) Type compatibility o Is language nominally-typed or structurally-typed? Parameter passing o How are params. passed (e.g. by value, by name)? Syntactic Features Name o Number of names which overlap Text o Number of text data which overlap Keyword o Number of keywords which overlap Literal o Number of literals which overlap Punctuation o Number of punctuation signs which overlap Operator o Number of operators which overlap Comment o Number of comment tokens which overlap Syntax o Number of AST nodes which overlap Tokens o Number of tokens which overlap Dataset-Specific Features Difficulty $x$ Average difficulty of dataset problems Length $x$ Average number of tokens Time limit $x$ Average time limit of dataset problems Memory limit $x$ Average memory limit of dataset problems Model-Specific Features Pretrained s Source language is included during pretraining Pretrained t Target language is included during pretraining Pretrained b Both source and target languages in pretraining
TABLE I: Features of language pairs. Column “Measure” indicates how the feature is computed: (e)quivalence, (o)verlap, (s)ource, (t)arget, (b)oth source and target, (rd) for relative difference. For dataset specific features, ($x$) can be s, t, or rd. Details of the linguistic features are available in supplementary material Tab. IX
Given a task and target language, how should we pick the source language for best performance? Knowing this without having to train many models on different source languages can save significant resources. It also provides a basis to study important features that characterize a successful transfer from a language to another. To predict the performance of the models on the different tasks for different source languages, we train a ranking model based on gradient-boosted decision trees (GBDT) [27]. We show that our models outperform baselines based on regressors or programming-language history. We consider the features defined in Tab. I. These features can be grouped into four categories.
  1. 1.
    Linguistic features: general properties characterizing a language pair like whether they support the same paradigms, the same style of memory management (is there a garbage collector or not), or the same type-system characteristics.
  2. 2.
    Syntactic features: properties of the syntax of a language pair measured by overlap in the counts of certain token types, as determined using the Pygments library on the corpus of code in the given languages.
  3. 3.
    Dataset-specific features: properties of the problems associated with code samples for the language pair in the online coding judge website from which the dataset came.
  4. 4.
    Model-specific features: whether the model saw the source language, target language, or both during pretraining.
To predict the top source languages for a given task and target language $L_{t}$ in the set of target languages $T$, we train a ranker using the LightGBM [27] implementation of the LambdaRank algorithm [28]. The model takes features in Tab. I as inputs and scores source languages $L_{s}$ in terms of their relevance to the target $L_{t}$. Our ranker uses a boosting ensemble of 100 decision trees with 16 leaves each. We consider the normalized discounted cumulative gain score (NDCG@K) at $K=3$ as our evaluation metric. We evaluate the model using leave-one-out (LOO) cross-validation on the set $T$. For each target language $L_{t}$, we train a ranker to predict rankings of different sources for each language in $T$ leaving out the source ranking for $L_{t}$ as a test set. For each fold, we compute the NDCG@3 score on the test set. We compare the performance of our ranker with a regression-based ranker and a history ranker. The history ranker selects the top $K$ source languages with the shortest path to the given target language in a programming languages history graph [^2] http://rigaux.org/language-study/diagram.html (contracted to merge different versions of the same language). The regression-based ranker uses a LightGBM regressor with the same hyperparameter settings as our ranker and selects the top $K$ source languages with the highest predicted scores. Fig. 6 shows the mean and standard deviation of NDCG@3 using LOO on $T$. While the error bars are overlapping, both regression and rankers outperform the history ranker on all tasks, indicating the importance of considering a variety of features for predicting top-performing sources rather than simplified heuristics. Our model outperforms baseline rankers on all tasks with the exception of clone detection, for which it shows slightly lower performance than the regression model. This indicates that ranking is the best suited method for predicting top performing sources. Based on these results, we use our ranker for the feature analysis in RQ3.
Figure 6: NDCG@3 scores for different rankers and tasks corresponding to LOO evaluation over the set of target languages.
Using a ranking model and considering a variety of features rather than simplified heuristics is crucial for predicting top performing sources.
IV-C Feature analysis
Figure 7: Normalized SHAP values aggregated by tasks for the features defined in Tab. I. The features are sorted by mean rank. Grayed-out features were unavailable for the given task.
While Sec. IV-A explored how well transfer learning performs for different languages on different tasks, we would like to dig deeper into why they perform that way. That is, how do the features of language pairs affect transfer? Which characteristics of the source and target language best explain the variation of transfer across tasks, the dependence of transfer on source and target pairing, as well as the superior transfer supported by specific sources? These questions can be answered by measuring feature importance.
Unfortunately, most traditional techniques for measuring feature importance suffer from a common pitfall: when two features overlap, they arbitrarily assign more importance to one at the expense of the other. The state-of-the-art technique for measuring feature importance avoids this pitfall by considering all possible subsets $S\subseteq F$ of features [29]. Then, the importance of a given feature is the difference it makes to model performance in combination with all possible sets of other features. The resulting SHAP values [29] show the relative impact of each feature on the model output. In the context of ranking different source languages for cross-lingual transfer, SHAP values measure how different features contribute to transfer. Fig. 7 shows SHAP values computed based on the ranking models evaluated in Sec. IV-B. For presentation purposes, it rescales the values so the most important feature for each task has importance 1.
Task-dependent feature importance
While previous works emphasize the importance of specific features for cross-lingual transfer, the lack of comparison on different tasks limits their conclusions. The supported cross-task analysis demonstrates a task-dependent feature importance whereby different features contribute differently to cross-lingual transfer depending on the task. For example, the top features for the solution domain classification task are Difficulty, Literals, Names, and Operators. Being able to predict the solution domain for a code solution requires some knowledge of the underlying problem for which the tags are an attribute. The difficulty score is another attribute of the problem. Problems with similar difficulty scores require similar algorithms. The literals, names, and operators are other indicators of the algorithms used in a code sample. In comparison, the most significant features for the clone detection task are Token, Names, and Keywords. Detecting a clone requires different skills than classifying domains. A deeper understanding of the code semantics irrespective of the similar algorithms used is needed. The overlaps in names, keywords, and more generally tokens is key for understanding the semantics of code [11].
Range of important features
Different tasks seem to not only focus on different features, but also focus on a different number of features. While the feature importance for solution domain classification and error detection is concentrated on a small number of features, the feature importance is more spread out for clone detection and code repair. For example, the code repair task requires a transfer of knowledge from overlapping keywords, names, and more generally different tokens from a source language for fixing a code in a transfer language. On the other hand, solution domain classification focuses on fewer features, related to the problems attributes.
Comparison across categories
Comparing the importance of different feature categories across tasks, we observe a relatively higher significance of syntactic features as demonstrated by the top two features Keywords and Names. Compared to transfer learning across natural languages [17], model-specific and linguistics features seem to carry less significance for transfer in programming languages. While de Vries et al. [17] find the potential of cross-lingual transfer dropping significantly for sources and targets that are not seen by the model, our finding brings hope to not only transferring learning to low-resource target languages, but even finetuning the model on unseen languages that could serve as good sources, such as Kotlin.
Within-category differences
While different categories seem to have a higher impact on the model predictions, the impact of single features within a category varies across task. Our results are consistent with prior work that found Names and Keywords to be the most important features for cross-lingual transfer in certain tasks [11]. However, half of the tasks show keywords to be more important than names, while the other half has the opposite order of importance.
1. The importance of language pair features for cross-lingual transfer performance depends on the task. 2. Higher-level tasks seem to draw upon a broader range of features to best understand the code. 3. Overall, the most important features across the four tasks studied in this paper are keywords and names.
V Threats to Validity
V-A External validity
One threat is that to make the large number of experiments feasible and reduce their carbon footprint, we had to limit ourselves to a moderately-sized model. We mitigate this threat by picking CodeT5 [22], which is widely used in the research literature for AI-for-Code. We believe that even if we had used larger models, insights from one model would not perfectly generalize to another.
Another threat is that to cover many languages, we restricted ourselves to four tasks. We mitigate this threat by picking tasks that cover a broad spectrum of difficulty and are diverse, as evidenced by their different performance and feature importances. Overall, our number of tasks exceeds prior work on cross-lingual transfer for programming languages.
A further threat is that our datasets have some synthetic aspects, such as type-IV clones [25] for clone detection data or fault injection for code repair data. However, using fully natural data is such a high standard that hardly any academic paper satisfies it, nor is it typically the case in industry practice. Instead, we mitigate this threat by explicitly measuring the importance of dataset-specific features of language pairs.
Another threat is that we focused on the setting of zero data in the target language. This represents the lowest barrier to entry, as gathering even a small number of examples can be tricky. Therefore, we leave exploring the case where there is some target-language data to future work.
V-B Internal validity
We believe that our study avoids one common internal validity threat, which revolves around insufficient runs. Thanks to the sheer number of language pairs for each task, we get thousands of different experiments, whose results are consistent with each other. Moreover, each of these experiments involves fine-tuning with tens of thousands of samples, giving us confidence that the choice of samples is not a threat either.
Since our study is the first of its kind at this scale, there is a threat that it cannot be directly compared to baselines. We mitigate this by observing that our results do confirm more narrow studies where applicable, for instance, a study of transfer for identifier-centric tasks [11].
VI Conclusion
We perform a systematic and extensive study of LLM transfer learning covering up to 41 programming languages, across 4 tasks including both classification and generation. To understand relative differences in cross-lingual performance between different programming languages, we define several linguistic, syntactic, dataset, and model-specific features of language pairs and then analyze the feature importance for a model that predicts transfer performance. Overall, we hope that this paper helps advance the community’s understanding of how learning transfers among programming languages, and that this improved understanding leads to better models to assist users of those languages, particularly low-resource ones.
VII Data Availability
The experiments are based on the publicly available CodeT5-base (220M parameters) model [22] and the open-sourced datasets CodeNet [19] and XCodeEval [20]. Upon paper acceptance, we plan to open-source the code to create the derived datasets and the framework to train and evaluate the models. Also, we plan to make the 58 finetuned models available with the final version of the paper.
We also plan to submit a replication package to the artifact committee.